A blower handles 1 kg/s of air at 20 degree cel and consumes a power of 15 kW. The inlet and outlet velocities of air are 100 m/s and 150 m/s respectively. Find the exit air temperature, assuming adiabatic conditions. Take cp of air i 1.005 kJ/kg K
1
Expert's answer
2021-10-28T08:53:01-0400

"c_p(t_2-t_1)=h_2-h_1,"

"w_1(h_1+\frac{v_1^2}{2000}+\frac{gz_1}{1000})=w_2(h_2+\frac{v_2^2}{2000}+\frac{gz_2}{1000})+W,"

"w_1=w_2,~z_1=z_2,"

"h_2-h_1=15+\frac{100^2}{2000}-\frac{150^2}{2000}=8.75,"

"t_2=20+\frac{8.75}{1.005}=28.7°C."